OpenStudy (anonymous):

How do you find a vertical asymptote in calculus? I know how to do it algebraically, but that is no good in Calc. Thanks in advance!

OpenStudy (zzr0ck3r):

why is it no good in calc?

OpenStudy (anonymous):

Our teacher wants us to use limits and stuff. Its just not how she wants us to do it

OpenStudy (zzr0ck3r):

well, you will get a asymptote at x, when the left and right limits are infinity and/or negative infinity at x.

OpenStudy (zzr0ck3r):

example \(\lim_{x\rightarrow 2}\frac{1}{x-2}=\infty\)

OpenStudy (anonymous):

But how do you know where to start? Do you have to do the algebra with the denominator first, and THEN use limits to prove that there are asymptotes at those values?

OpenStudy (zzr0ck3r):

yeah

OpenStudy (zzr0ck3r):

you need to know where you are looking first.

OpenStudy (anonymous):

Ok, I just wanted to know if you had to do some calculus thing to know where to look first, but that makes sense. Thanks a lot!!!

OpenStudy (zzr0ck3r):

np